For the past records about40% of a firm orders are for export. Their record for export is 48% in one particular financial quarter. If they expect to satisfy about 80 orders in the next financial quarter, what is the probability that they will break their previous export record
1
Expert's answer
2020-03-13T09:59:02-0400

Normal approximation of the binomial distribution with

"\mu=np=80*0.48=38.4,\;\;\sigma=\sqrt{np(1-p)}=\sqrt{80*0.48(1-0.48)}\approx4.47"

can be used.

So, the probability that they will break their previous export record "80*0.4=32" orders is

"P(X>32)=P(Z>\frac{32-38.4}{4.47})=P(Z>-1.43)=0.9236."